# Seeded Poisson Factorization: Leveraging domain knowledge to fit topic models

### v1

@@ Abstract
Topic models are widely used for discovering latent thematic structures in large text corpora, yet traditional unsupervised methods often struggle to align with predefined conceptual domains. This paper introduces Seeded Poisson Factorization (SPF), a novel approach that extends the Poisson Factorization framework by incorporating domain knowledge through seed words. SPF enables a more interpretable and structured topic discovery by modifying the prior distribution of topic-specific term intensities, assigning higher initial rates to predefined seed words. The model is estimated using variational inference with stochastic gradient optimization, ensuring scalability to large datasets.
We apply SPF to an Amazon customer feedback dataset, leveraging predefined product categories as guiding structures. Our evaluation demonstrates that SPF achieves superior classification performance compared to alternative guided topic models, particularly in terms of computational efficiency and predictive performance. Furthermore, robustness checks highlight SPF’s ability to adaptively balance domain knowledge and data-driven topic discovery, even in cases of imperfect seed word selection. These results establish SPF as a powerful and scalable alternative for integrating expert knowledge into topic modeling, enhancing both interpretability and efficiency in real-world applications.
@@ 1 Introduction
Inferring latent structures in text data is a fundamental challenge in natural language processing and its application in a wide range of fields of research such as political science, social science and economics. Due to the unstructured nature of text data, text analysis poses distinct challenges compared to the analysis of other types of data that are commonly used in empirical research (see, e.g., Kelly et al.,, 2021). Topic modeling provides a widely used framework for discovering hidden thematic structures within text corpora, offering insights into the distribution of topics across documents and the association between words and topics. Among the available topic modeling approaches, in particular Latent Dirichlet Allocation (LDA; Blei et al.,, 2003) and its extensions (see, e.g., Lafferty and Blei,, 2005; Roberts et al.,, 2014; Eshima et al.,, 2024) have been widely studied and applied across various domains (see, e.g., Bagozzi and Berliner,, 2018; Barbera et al.,, 2019; Thorsrud,, 2020; Davis and Tabrizi,, 2021; Munro and Ng,, 2022; Zimmermann et al.,, 2024). However, alternative topic modeling frameworks, such as Poisson Factorization (PF), provide distinct advantages by leveraging a Poisson likelihood rather than a multinomial distribution.
@@
PF has been shown to provide improved scalability and computational efficiency (see, e.g., Canny,, 2004; Gopalan et al.,, 2014, 2015). Unlike LDA, which relies on Dirichlet priors for topic distributions, PF leverages the Poisson distribution. PF factorizes the document-term matrices into non-negative latent components that naturally accommodate sparsity, making it well suited for large-scale text corpora (Gopalan et al.,, 2015). In the case of PF topic models, these latent components represent topic intensities over words $\bm{\beta}$ (referred to as topical content or topic-term intensities), and document intensities over topics $\bm{\theta}$ (referred to as topical prevalence or document-topic intensities). The topical content refers to what is being discussed, while the topical prevalence indicates how much it is being discussed. PF naturally promotes sparsity in the latent factors and can handle large datasets efficiently due to its inherent properties and the use of variational inference techniques (see, e.g., Vafa et al.,, 2020; Vávra et al., 2024a, ; Vávra et al., 2024b, ).
Despite the success of topic models in uncovering latent themes in textual data, traditional methods are often limited by their purely unsupervised nature. In many applications, researchers and practitioners require models that align with predefined conceptual domains or that allow for targeted analysis based on domain knowledge (see, e.g., Eshima et al.,, 2024, for an application in political science). Conventional topic models typically require extensive post-processing, such as the manual labeling of topics and the merging of semantically similar ones, to extract meaningful insights (see, e.g., Morstatter and Liu,, 2016; Eshima et al.,, 2024). To address this limitation, guided and semi-supervised topic models have been proposed, which incorporate human input a-priori to fit a topic model to learn topics of specific interest to a user (see, e.g., Jagarlamudi et al.,, 2012; Watanabe and Zhou,, 2022; Eshima et al.,, 2024). These approaches typically influence the topic-term distributions, allowing for more interpretable and controlled topic discovery. Recent advances in guided topic modeling have primarily focused on extending LDA-based models (Watanabe and Zhou,, 2022; Eshima et al.,, 2024; Watanabe and Baturo,, 2024), yet a corresponding extension within the PF framework has remained unexplored.
@@
The paper at hand contributes to the literature as follows. Firstly, we introduce Seeded Poisson Factorization (SPF), a novel topic model that integrates domain knowledge into the PF framework through the inclusion of seed words. As shown in Figure 1, SPF extends standard PF by decomposing topic-term intensities into a neutral component and a seeded component informed by predefined seed words. This structured decomposition enables the model to incorporate prior knowledge while preserving the flexibility of PF to learn latent topic distributions from data. In this way, SPF learns specific topics of interest and adaptively adjusts for potential seed word mis-specifications by controlling the contribution of seeded components, ensuring robustness in applications where domain knowledge may be incomplete or imperfect. Secondly, in contrast to prior guided topic models, which predominantly rely on Markov Chain Monte Carlo (MCMC) inference (see, e.g., Eshima et al.,, 2024; Watanabe and Baturo,, 2024), SPF employs variational inference (VI) for scalable parameter estimation. VI formulates posterior inference as an optimization problem, significantly reducing computational costs compared to traditional sampling-based methods (Ranganath et al.,, 2014; Blei et al.,, 2017). We empirically demonstrate that SPF not only achieves competitive predictive performance but also exhibits substantial computational efficiency, making it particularly suitable for large-scale text corpora.
@@
Topic models are used for automatic text analysis in a wide range of fields of research, including for example the analysis of speech data in political science (Vávra et al., 2024b, ) or open-ended survey responses in the social sciences (Roberts et al.,, 2014). In the following we focus on yet another area of application: automatic analysis of consumer feedback. Consumer feedback provides valuable insights into customer preferences, sentiment, and emerging trends (see, e.g., Khan and Jeong,, 2016; Filieri et al.,, 2018; Davis and Tabrizi,, 2021; Aghakhani et al.,, 2021; Biswas et al.,, 2022; Zhou et al.,, 2024). Given the growing volume of online reviews and their impact on decision-making, accurately categorizing and summarizing such feedback remains a central challenge in computational social science and business analytics. We thus make use of a publicly available Amazon customer review dataset to illustrate and evaluate the performance of SPF. In particular, we assess SPF’s performance to extract meaningful topics when seed words are supplied to characterize the underlying product categories of products discussed in the customer reviews. In addition we investigate how well SPF infers the product category of the product discussed in a consumer review by employing a Naive Bayes classifier on the inferred topic intensities. We also compare the predictive performance as well as the computational efficiency of SPF to competing recently proposed guided topic models with readily available software implementations, i.e., KeyATM (Eshima et al.,, 2024) and SeededLDA (Watanabe and Baturo,, 2024). Finally, we conduct a series of robustness checks to examine the sensitivity of SPF to variations in seed word quality, model specification and corpus characteristics. The results confirm that SPF provides excellent performance across diverse experimental conditions.
@@
By introducing SPF, we contribute to the methodological literature on topic modeling by extending Poisson Factorization with domain-informed priors, providing an alternative to existing LDA-based guided topic models. Our results demonstrate that SPF enhances both interpretability and computational efficiency of topic models, offering a scalable solution for researchers and practitioners seeking structured topic discovery in large text corpora.
@@
The rest of the paper is structured as follows. In Section 2 we describe our generative model. Section 3 outlines model inference. Section 4 provides empirical results by applying SPF to customer feedback data from Amazon. Section 5 concludes.
@@ 2 The Seeded Poisson Factorization model
Based on the bag-of-words assumption (see, e.g., Eshima et al.,, 2024) the data are summarized in a Document-Term-Matrix (DTM), $\mathbb{Y}$. This matrix has the dimension number of documents $D$ times number of unique terms (words) in the data $V$, where each row corresponds to a single document $d=1,\ldots,D$, and each column represents a specific term $v=1,\ldots,V$ from vocabulary $\mathscr{V}$. A single entry $y_{dv}$ contains the frequency count of term $v$ in document $d$, such that $y_{dv}\geq 0$. Poisson factorization topic models assume that the observed word frequencies are generated independently from a Poisson distribution. In our Seeded Poisson Factorization (SPF) model, we decompose the Poisson rates into a linear combination of document-topic intensities $\bm{\theta}$ and topic-term intensities $\bm{\beta}$ over latent topic dimension $K$ for every frequency count $y_{dv}$:
$y_{dv}\sim\mathsf{Pois}\left(\sum_{k=1}^{K}\theta_{dk}\beta_{kv}\right).$ (1)
@@
In our SPF model we include the prior knowledge available to researchers which suggests specific terms to be relevant for certain topics in the following way. We seed the topics by inflating the prior mean of topic-term intensities for seeded words. Let $\mathscr{V}_{k}\subset\mathscr{V}$ be the set of seed words for topic $k=1,\ldots,K$ of size $V_{k}=|\mathscr{V}_{k}|$. In practice, we expect only few seed words per topic, $V_{k}\ll V$ and denote by $\mathscr{S}=\bigcup_{k=1}^{K}\mathscr{S}_{k}$, $\mathscr{S}_{k}=\{(k,v),v\in\mathscr{V}_{k}\}$, the set of all seed words. We allow for $\mathscr{V}_{k}=\emptyset$, $V_{k}=0$, in which case the topic is not a-priori seeded. Note that in case $V_{k}=0$ for all $k=1,\ldots,K$, the SPF model reduces to a standard PF model.
@@
with $a,b,c,d>0$ and $c\gg a$. In case $b=d$, this implies that $\beta_{kv}$ also has a gamma prior $\Gamma(a+c,b)$. All document-topic intensities $\theta_{dk}$ are given a gamma prior
@@
To obtain the empirical results in Section 4, we set $a=b=e=f=0.3$, which according to Gopalan et al., (2015) results in sparse representations of the document-topic and topic-term intensities. To emphasize the relevance of the pre-defined seed words we set $c=1.0$ and $d=0.3$. Finally, the generative process is captured in plate notation in Figure 2.
@@ 3.1 Variational inference
Given the DTM $\mathbb{Y}$, we infer the document-topic intensities and the topic-term intensities based on approximating the posterior distribution over the model’s latent variables $p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}}\,|\,\mathbb{Y})$. We use Variational Inference (VI) methods to fit an approximate posterior distribution (see, e.g., Blei et al.,, 2017). VI frames the inference as an optimization problem. The key steps of VI consist of selecting a parametric family of variational distributions $\mathscr{Q}=\{q_{\bm{\phi}},\bm{\phi}\in\bm{\Phi}\}$ and determining the parameter $\bm{\phi}^{*}\in\bm{\Phi}$ minimizing the Kullback-Leibler divergence ($\mathsf{KL}$) of the variational distribution from the true posterior
$q_{\bm{\phi}^{*}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})=\underset{q_{\bm{\phi}}\in\mathscr{Q}}{\operatorname{\mathrm{arg}\,\mathrm{min}\;}}\mathsf{KL}\left(q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})\,\left\|\,p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}}\,|\,\mathbb{Y})\right.\right).$
@@
$\mathsf{ELBO}(\bm{\phi})=\mathbb{E}_{{q}_{\bm{\phi}}}\left[\log p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})+\log p(\mathbb{Y}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})-\log q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})\right]$ (4)
or minimizing the negative $\mathsf{ELBO}$ (Jordan et al.,, 1998; Bishop,, 2006). Equation 4 sums the expectation of the log-likelihood and the log-prior and the entropy of the variational family.
@@
$q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$ $=\prod_{d,k=1}^{D,K}q(\theta_{dk})\prod_{k,v=1}^{K,V}q(\beta^{\star}_{kv})\prod_{k,v\in\mathscr{S}}q(\widetilde{\beta}_{kv}).$
Only distributions with support on the positive reals are suitable as variational distributions for document-topic intensities and topic-term intensities. We propose to use Gamma distributions, matching the prior distributions outlined in (2) and (3). We denote shape and rate parameters with the superscript ‘$\mathsf{shp}$’ and ‘$\mathsf{rte}$’, respectively. Moreover, we also employ a scaling by document length $N_{d}=\sum\limits_{v=1}^{V}y_{dv}$ for parameters $\theta_{dk}$. Including the document length $N_{d}$ in this way provided a more stable and quicker model fit and induced a superior classification performance. Altogether, we posit as variational distributions
@@
We use Black Box Variational Inference (BBVI) with stochastic optimization and follow Ranganath et al., (2014) to form noisy unbiased gradient estimates of the $\mathsf{ELBO}$ with $S$ Monte Carlo samples from the variational distribution,
@@
where $\bm{\theta}_{s},\bm{\beta^{\star}}_{s},\widetilde{\bm{\beta}}_{s}\sim q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$ are independent samples from the variational distributions. These gradient estimates are used to optimize the $\mathsf{ELBO}$ while the updates $\bm{\phi}$ are determined by the Adam algorithm (Kingma and Ba,, 2015). Reverse-mode automatic differentiation is used to track all sequences of operations and to compute the gradients during the optimization procedure (see Kucukelbir et al.,, 2017). The whole procedure is shown in Algorithm 1 for $S=1$ which is the value for $S$ we use in our implementation. When applying Algorithm 1, we initialize the variational parameter with $\bm{\phi}^{\mathsf{shp}}_{\theta}=1$, $\bm{\phi}^{\mathsf{rte}}_{\theta}=\frac{D}{1000}$, $\bm{\phi}^{\mathsf{shp}}_{\beta^{\star}}=1$, $\bm{\phi}^{\mathsf{rte}}_{\beta^{\star}}=\frac{2D}{1000}$, $\bm{\phi}^{\mathsf{shp}}_{\widetilde{\beta}}=\bm{\phi}^{\mathsf{rte}}_{\widetilde{\beta}}=1$.
@@
Input: number of epochs $E$, batch size $|\mathscr{B}|$, learning rate $\rho$
Output: The last value ^$\bm{\hat{}}{\phi}$ when optimizing $\mathsf{ELBO}(\bm{\phi})$
@@
3       Divide $D$ documents randomly into $B$ batches $\mathscr{B}_{i}$, $i=1,\ldots,B$, $|\mathscr{B}_{i}|\approx|\mathscr{B}|$
4      for i in $1:B$ do
@@
6                   Sample $\beta^{\star}_{kv}\sim\Gamma(\phi^{\mathsf{shp}}_{\beta^{\star}_{kv}},\phi^{\mathsf{rte}}_{\beta^{\star}_{kv}})$
@@
8                         Sample $\widetilde{\beta}_{kv}\sim\Gamma(\phi^{\mathsf{shp}}_{\widetilde{\beta}_{kv}},\phi^{\mathsf{rte}}_{\widetilde{\beta}_{kv}})$
@@ 10                         Set $\widetilde{\beta}_{kv}=0$
12            Compute $\bm{\beta}=\bm{\beta^{\star}}+\bm{\widetilde{\beta}}$
@@
14            for each document $d$ in batch $\mathscr{B}_{i}$ do
15                   Sample $\theta_{dk}\sim\Gamma(\phi^{\mathsf{shp}}_{\theta_{dk}},N_{d}\cdot\phi^{\mathsf{rte}}_{\theta_{dk}})$
@@
18                         Set $\lambda_{dv}=\sum\limits_{k=1}^{K}\theta_{dk}\beta_{kv}$
                         Compute $\log p(y_{dv}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})=\log\mathsf{Pois}(y_{dv}\,|\,\lambda_{dv})$
@@ 18                         Set $\lambda_{dv}=\sum\limits_{k=1}^{K}\theta_{dk}\beta_{kv}$
            Set $\log p(\mathbb{Y}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})=\frac{D}{|\mathscr{B}_{i}|}\sum\limits_{d\in\mathscr{B}_{i}}\sum\limits_{v=1}^{V}\log p(y_{dv}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$
@@
            Compute $\log p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$ and $\log q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$
@@
22             Compute $\mathsf{ELBO}(\bm{\phi})=\log p(\mathbb{Y}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})+\log p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})-\log q_{\bm{\phi}}(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})$
23             Compute gradients $\nabla_{\phi}\mathsf{ELBO}(\bm{\phi})$ using automatic differentiation as in (5)
24             Update variational parameter $\bm{\phi}$ with Adam and learning rate $\rho$
@@
3.2 Post-processing and final inference
After running the model for a sufficient number of epochs $E$, the final value obtained from the VI optimization $\hat{\bm{\phi}}$ represents the estimate of the variational parameter. We summarize the results by determining point estimates for the parameters of interest based on posterior means derived from the posterior approximations through the variational family. In particular, the posterior mean estimates $\hat{\bm{\theta}},\hat{\bm{\beta}}^{\star},\hat{\widetilde{\bm{\beta}}}$ are obtained by determining the means induced by the variational Gamma distributions. In case of document-topic intensities we obtain posterior means using

### v2

@@ Abstract
Topic models are widely used for discovering latent thematic structures in large text corpora, yet traditional unsupervised methods often struggle to align with pre-defined conceptual domains. This paper introduces seeded Poisson Factorization (SPF), a novel approach that extends the Poisson Factorization (PF) framework by incorporating domain knowledge through seed words. SPF enables a structured topic discovery by modifying the prior distribution of topic-specific term intensities, assigning higher initial rates to pre-defined seed words. The model is estimated using variational inference with stochastic gradient optimization, ensuring scalability to large datasets.
We present in detail the results of applying SPF to an Amazon customer feedback dataset, leveraging pre-defined product categories as guiding structures. SPF achieves superior performance compared to alternative guided probabilistic topic models in terms of computational efficiency and classification performance. Robustness checks highlight SPF’s ability to adaptively balance domain knowledge and data-driven topic discovery, even in case of imperfect seed word selection. Further applications of SPF to four additional benchmark datasets, where the corpus varies in size and the number of topics differs, demonstrate its general superior classification performance compared to the unseeded PF model.
@@ 1 Introduction
Inferring latent structures in text data is a fundamental challenge in natural language processing and its application in a wide range of fields of research such as political science, social science and economics. Due to the unstructured nature of text data, text analysis poses distinct challenges compared to the analysis of other types of data that are commonly used in empirical research [87, see, e.g.,]. Topic modeling provides a widely used framework for discovering hidden thematic structures within text corpora, offering insights into the distribution of topics across documents and the association between words and topics. Among the available topic modeling approaches, in particular Latent Dirichlet Allocation [69, LDA;] and its extensions [92, 107, 76, see, e.g.,], which use the document-term matrix as input, have been widely studied and applied across various domains [64, 65, 110, 73, 102, 122, 98, 72, see, e.g.,]. However, alternative topic modeling frameworks, such as Poisson Factorization (PF), provide distinct advantages by leveraging a Poisson likelihood rather than a multinomial distribution and providing a more flexible prior parameter specification compared to LDA.
@@
PF has been shown to provide a better fit to the data as well as improved scalability and computational efficiency [70, 79, 80, see, e.g.,]. PF factorizes the document-term matrices into non-negative latent components, which correspond to topic intensities over words $\bm{\beta}$ (referred to as topical content or topic-term intensities), and document intensities over topics $\bm{\theta}$ (referred to as topical prevalence or document-topic intensities). The topical content refers to what is being discussed, while the topical prevalence indicates how much it is being discussed. PF naturally promotes sparsity and can handle large datasets efficiently due to its inherent properties and the use of variational inference techniques [112, 83, 113, 114, see, e.g.,].
Despite the success of topic models in uncovering latent themes in textual data, traditional methods are often limited by their purely unsupervised nature. In many applications, researchers and practitioners require models that align with pre-defined conceptual domains or that allow for targeted analysis based on domain knowledge [76, see, e.g.,]. Extensions of topic models to allow for guidance, e.g., via the inclusion of seed words, have thus been considered in a number of contributions, indicating their suitability to improve interpretability of topics as well as the use for automatic text classification. In this context, one can differentiate in particular between approaches extending non-probabilistic topic models, approaches extending the LDA-based probabilistic topic model and methods to improve the creation of seed words.
The stream of literature extending non-probabilistic topic models includes among others [78] who pursue in their proposed anchored CorEx algorithm an information-theoretic framework which enforces a single-membership of words within topics and takes seed words into account as anchor words using as input a binarized version of the document-term matrix. Furthermore, exploiting the recent advances in large language models, [104] propose TopicGPT to uncover latent topics in a text collection based on a concise label paired with a broad one-sentence description to characterize a topic by prompting these models. [81] combines in BERTopic document embeddings generated with pre-trained transformer-based language models with clustering of these embeddings and generating topic representations with a class-based term-frequency-inverse document frequency procedure, where an extension also allows for an inclusion of seed words.
LDA-based extensions of the probabilistic topic model to guide the topic estimation are considered by a number of contributions including [84, 96, 95, 82, 97, 116, 76, 115]. These approaches typically influence the topic-term distributions, thereby enhancing interpretability and enabling more controlled modeling.
To improve guiding the estimation of topic models, approaches to derive a suitable set of seed words have also been investigated. These methods may be employed to obtain the seed words used as input to the guided topic model approaches and thus improve the overall performance. In this context, [99] propose the CatE approach, which learns a discriminative embedding space and discovers category representative terms in an iterative manner based on category names. [120] propose their iterative framework SeedTopicMine, which allows them to jointly learn from three types of context suitable sets of terms to be used as seed words.
Probabilistic topic models are built either on LDA or on PF. Although extensions of the LDA model to include guidance and seed words are numerous, we are not aware of work that extended PF in this direction, despite other extensions of the PF framework considered in the literature, such as for example [75] where the inclusion of a topic hierarchy is considered. The paper at hand thus contributes to this literature by introducing a topic model using seed words within the PF framework.
@@
In particular, this paper contributes to the literature as follows. Firstly, we introduce seeded Poisson Factorization (SPF), a novel topic model that integrates domain knowledge into the PF framework through the inclusion of seed words. As shown in Figure 1, SPF extends standard PF by decomposing topic-term intensities into a neutral component and a seeded component informed by pre-defined seed words. This structured decomposition enables the model to incorporate prior knowledge while preserving the flexibility of PF to learn latent topics from data. In this way, SPF learns specific topics of interest and adaptively adjusts for potential seed word misspecifications by controlling the contribution of seeded components, ensuring robustness in applications where domain knowledge may be incomplete or imperfect. Secondly, in contrast to prior guided topic models, which predominantly rely on Markov Chain Monte Carlo (MCMC) inference [76, 115, see, e.g.,], SPF employs variational inference (VI) for scalable parameter estimation. While MCMC methods are in principle applicable, we rely on VI methods, which formulate posterior inference as an optimization problem, significantly reducing computational costs compared to traditional sampling-based methods [105, 68]. We empirically demonstrate that SPF not only achieves competitive predictive performance but also exhibits substantial computational efficiency, making it particularly suitable for large-scale text corpora.
@@
Topic models are used for automatic text analysis in a wide range of fields of research, including for example text classification in data journalism [109], open-ended survey responses in the social sciences [107] and analysis of speech data in political science [114]. In the following, we focus on yet another area of application: automatic analysis and classification of consumer feedback. Consumer feedback provides valuable insights into customer preferences, sentiment, and emerging trends [88, 63, 77, 73, 62, 119, 67, 121, see, e.g.,]. Given the growing volume of online reviews and their impact on decision-making, accurately categorizing and summarizing this feedback remains a central challenge in computational social science and business analytics. Thus, we make use of a publicly available Amazon customer review dataset to illustrate and evaluate the performance of SPF. In particular, we assess SPF’s performance to extract meaningful topics when seed words are supplied to characterize the underlying product categories of products discussed in the customer reviews. In addition, we investigate how well SPF infers the product category of the product discussed in a consumer review by employing a Naive Bayes classifier on the inferred topic intensities. We also compare the predictive performance as well as the computational efficiency of SPF to competing recently proposed guided topic models with readily available software implementations, i.e., KeyATM [76] and SeededLDA [115]. We conduct a series of robustness checks to examine the sensitivity of SPF to variations in seed word quality, model specification and corpus characteristics. In addition, we also fit SPF to four publicly available corpora with known categories to indicate the general applicability of our model for automatic text classification. Our results confirm that SPF provides excellent performance aunder various experimental conditions.
@@
By introducing SPF, we contribute to the methodological literature on topic modeling by extending PF with domain-informed priors, providing an alternative to existing LDA-based guided topic models. Our results demonstrate that SPF enhances both the guidance and computational efficiency of topic models, offering a scalable solution for researchers and practitioners seeking structured topic discovery in large text corpora.
@@
The rest of the paper is structured as follows. In Section 2, we describe our generative model. Section 3 outlines the model inference. Section 4 presents empirical results based on the application of SPF to Amazon customer feedback data as well as four benchmark datasets consisting of text corpora and their categorization. Section 5 concludes.
@@ 2 The seeded Poisson factorization model
Based on the bag-of-words assumption [76, see, e.g.,] the data are summarized in a Document-Term-Matrix (DTM), $\mathbb{Y}$. This matrix has the dimension number of documents $D$ times number of unique terms (words) in the data $V$, where each row corresponds to a single document $d=1,\ldots,D$, and each column represents a specific term $v=1,\ldots,V$ from vocabulary $\mathscr{V}$. A single entry $y_{dv}$ contains the frequency count of term $v$ in document $d$, such that $y_{dv}\geq 0$. PF topic models assume that the observed word frequencies are generated independently from a Poisson distribution. The Poisson rates are decomposed into a linear combination of document-topic intensities $\bm{\theta}$ and topic-term intensities $\bm{\beta}$ over latent topic dimension $K$ for every frequency count $y_{dv}$:
$y_{dv}\sim\mathsf{Pois}\left(\sum_{k=1}^{K}\theta_{dk}\beta_{kv}\right).$
@@
In its standard form, the PF models the frequency of words in documents without including any prior knowledge about the topic structure. However, in many applications, researchers possess domain knowledge that suggests certain terms that are highly indicative of specific topics. To leverage this information, we extend the standard PF by introducing structured priors based on the inclusion of seed words. Specifically, SPF differs from PF in that the topic-term intensities are decomposed into two components: a neutral component representing unsupervised topic discovery and a seeded component that emphasizes pre-identified important terms for specific topics. This structured decomposition allows the model to prioritize seed words during inference, steering the learned topics towards meaningful, interpretable structures aligned with user-specified domain knowledge.
In particular, we include the prior knowledge in SPF in the following way. We seed the topics by inflating the prior mean of topic-term intensities for seeded words. Let $\mathscr{V}_{k}\subset\mathscr{V}$ be the set of seed words for topic $k=1,\ldots,K$ of size $V_{k}=|\mathscr{V}_{k}|$. In practice, we expect only a few seed words per topic, $V_{k}\ll V$ and denote by $\mathscr{S}=\bigcup_{k=1}^{K}\mathscr{S}_{k}$, $\mathscr{S}_{k}=\{(k,v),v\in\mathscr{V}_{k}\}$, the set of all seed words. We allow for $\mathscr{V}_{k}=\emptyset$, $V_{k}=0$, in which case the topic is not a-priori seeded. Note that in case $V_{k}=0$ for all $k=1,\ldots,K$, the SPF model reduces to a standard PF model.
@@
with $a,b,c,d>0$ and $c\gg a$. In case $b=d$, this implies that $\beta_{kv}$ also has a gamma prior $\Gamma(a+c,b)$ if $(k,v)\in\mathscr{S}$. All document-topic intensities $\theta_{dk}$ are given a gamma prior
@@
To obtain the empirical results in Section 4, we set $a=b=e=f=0.3$, which according to [80] results in sparse representations of the document-topic and topic-term intensities. To emphasize the relevance of the pre-defined seed words, we set $c=1.0$ and $d=0.3$. Finally, the generative process is captured in plate notation in Figure 2.
@@ 3.1 Variational inference
Given the DTM $\mathbb{Y}$, we infer the document-topic intensities and the topic-term intensities based on approximating the posterior distribution over the model’s latent variables $p(\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}}\,|\,\mathbb{Y})$. We use Variational Inference (VI) methods to fit an approximate posterior distribution [68, see, e.g.,]. VI frames the inference as an optimization problem. The key steps of VI consist of selecting a parametric family of variational distributions $\mathscr{Q}=\{q_{\bm{\phi}},\bm{\phi}\in\bm{\Phi}\}$ and determining the parameter $\bm{\phi}^{*}\in\bm{\Phi}$ minimizing the Kullback-Leibler divergence ($\mathsf{KL}$) of the variational distribution from the true posterior
[…]
@@
[…]
or minimizing the negative $\mathsf{ELBO}$ [85, 66]. Equation (3) sums the expectation of the log-likelihood and the log-prior and the entropy of the variational family.
@@
[…]
Only distributions with support on the positive reals are suitable as variational distributions for document-topic and topic-term intensities. We propose to use gamma distributions, matching the prior distributions outlined in Equations (1) and (2). We denote shape and rate parameters with the superscript $\mathsf{shp}$’ and $\mathsf{rte}$’, respectively. Moreover, we also employ a scaling by document length $N_{d}=\sum\limits_{v=1}^{V}y_{dv}$ for parameters $\theta_{dk}$. Including the document length $N_{d}$ in this way provided empirically a more stable and quicker model fit and induced a superior classification performance. Altogether, we posit as variational distributions
@@
We use Black Box Variational Inference (BBVI) with stochastic optimization and follow [105] to form noisy unbiased gradient estimates of the $\mathsf{ELBO}$ with $S$ Monte Carlo samples from the variational distribution,
@@
[…]
@@
Input: number of epochs $E$, batch size $|\mathscr{B}|$, learning rate $\rho$.
Output: The last value $\bm{\hat{\phi}}$ when optimizing $\mathsf{ELBO}(\bm{\phi})$.
@@
3  Divide $D$ documents randomly into $B$ batches $\mathscr{B}_{b}$, $b=1,\ldots,B$, $|\mathscr{B}_{b}|\approx|\mathscr{B}|$.
4 for b in $1:B$ do
@@
6        Sample $\beta^{\star}_{kv}\sim\Gamma(\phi^{\mathsf{shp}}_{\beta^{\star}_{kv}},\phi^{\mathsf{rte}}_{\beta^{\star}_{kv}})$.
@@
8           Sample $\widetilde{\beta}_{kv}\sim\Gamma(\phi^{\mathsf{shp}}_{\widetilde{\beta}_{kv}},\phi^{\mathsf{rte}}_{\widetilde{\beta}_{kv}})$.
@@
12    Compute $\bm{\beta}=\bm{\beta^{\star}}+\bm{\widetilde{\beta}}$.
@@
14    for each document $d$ in batch $\mathscr{B}_{b}$ do
15        Sample $\theta_{dk}\sim\Gamma(\phi^{\mathsf{shp}}_{\theta_{dk}},N_{d}\cdot\phi^{\mathsf{rte}}_{\theta_{dk}})$.
@@
18           Set $\lambda_{dv}=\sum\limits_{k=1}^{K}\theta_{dk}\beta_{kv}$.
           Compute $\log p(y_{dv}\,|\,\bm{\theta},\bm{\beta}^{\star},\widetilde{\bm{\beta}})=\log\mathsf{Pois}(y_{dv}\,|\,\lambda_{dv})$.
@@
[…]
@@
[…]
@@
[…]
23     Compute gradients $\nabla_{\phi}\mathsf{ELBO}(\bm{\phi})$ using automatic differentiation as in Equation (4).
24     Update variational parameter $\bm{\phi}$ with Adam and learning rate $\rho$.
@@
3.2 Post-processing final inference
After running the model for a sufficient number of epochs $E$, the final value obtained from the VI optimization $\hat{\bm{\phi}}$ represents the estimate of the variational parameter. We summarize the results by determining point estimates for the parameters of interest based on posterior means derived from the posterior approximations through the variational family. In particular, the posterior mean estimates $\hat{\bm{\theta}},\hat{\bm{\beta}}^{\star},\hat{\widetilde{\bm{\beta}}}$ are obtained by determining the means induced by the variational Gamma distributions. In case of document-topic intensities, we obtain posterior means using